### v1

@@ 6 Discussion
The robustness of TabulaTime was tested under different environmental and temporal conditions. Notably, during the COVID-19 lockdown (2020–2021)—when air pollution levels were drastically reduced, our model maintains high predictive accuracy (78.8%), despite the altered exposure conditions. This result underscores TabulaTime’s adaptability to fluctuating environmental factors, making it potentially valuable for real-time health monitoring systems.
Clinically, TabulaTime’s interpretability offers actionable insights. The use of attention mechanisms and feature importance analysis enhances the interpretability of the model. Our analysis identified Systolic Blood Pressure (SBP), Body Mass Index (BMI), PM10, and NO2 as the most influential features for ACS risk prediction (Figure 9) . The high importance of PM10 and NO2 suggests a strong link between air pollution and acute cardiovascular events, reinforcing the need for air quality monitoring as a preventive healthcare measure. Additionally, the day-wise analysis of air pollution exposure (Figure 10) revealed that pollution levels one day prior to ACS onset were the most predictive, suggesting that short-term exposure triggers acute cardiovascular events. This result is clinically significant, as it supports policies advocating for real-time air quality alerts and public health interventions in high-pollution areas.
Despite its strengths, the TabulaTime model has several limitations. The model was tested using data from Salford, UK, so it may not work exactly the same way in other regions where air pollution levels and healthcare conditions are different. Future studies should test the model in other or larger locations to see if it can be applied more widely. Also, while PM10, and NO2 were identified as important pollutants, other environmental factors like temperature, humidity, and noise levels were not included in this study. Adding these in the future could make the predictions even more accurate.
In conclusion, TabulaTime is a powerful tool for ACS prediction that integrates environmental and clinical data. It is more accurate, efficient, and adaptable than traditional models, offering valuable insights for both clinical decision-making and public health policies. With further validation and refinement, it could play a significant role in personalized healthcare, air pollution monitoring, and real-time risk assessment.
7 Conclusion and Future work
This study introduces TabulaTime, a novel multimodal model that integrates time-series air pollution data with traditional clinical data to investigate their influence on Acute Coronary Syndrome (ACS) presentation. The results demonstrated that TabulaTime effectively incorporated these elements, achieving outstanding performance. Notably, the PatchRWKV module significantly enhanced the model’s accuracy by capturing both short-term and long-term pollution exposure trends, a major advancement in temporal feature extraction.
The experimental results underscored TabulaTime’s effectiveness in improving predictive accuracy across various tasks. Our TabulaTime outperformed other state-of-the-art machine learning models, such as Random Forests, LightGBM, and CatBoost, by margins ranging from 16.7% to 32.2%, showcasing its robustness and efficacy. Specifically, our findings highlighted the critical role of environmental factors in cardiovascular risk assessment. The inclusion of air quality data led to a 10.1% improvement in accuracy compared to scenarios where such data was omitted, emphasizing the importance of integrating environmental factors in ACS risk prediction—an aspect often overlooked in traditional models.
Furthermore, the PatchRWKV module demonstrated superior performance in time-series classification and forecasting tasks. In classification tasks, it outperformed state-of-the-art models such as LightTS, DLinear, TimesNet, Autoformer and PatchTST. In forecasting tasks, PatchRWKV consistently delivered lower Mean Squared Error (MSE) and Mean Absolute Error (MAE) compared to other models, demonstrating robust predictive capabilities for both weather and air pollution indicators.
The model maintained high accuracy despite significant fluctuations in air pollution levels, with accuracy rates of 82.9%, 78.8%, and 80.2% for pre-COVID, during COVID lockdown, and post-COVID periods, respectively. This adaptability suggested that TabulaTime was applicable across diverse geographical and temporal contexts. The feature importance analysis revealed that clinical factors such as previous angina and systolic blood pressure, alongside environmental pollutants like PM10 and NO2, are significant predictors of ACS risk. Interestingly, metabolic factors such as BMI and hypertension, while still relevant, were less influential in this model. This finding suggested that while metabolic factors contributed to cardiovascular risk, the interaction between clinical history and environmental exposure played a more pivotal role in the context of ACS prediction.
Overall, TabulaTime represents a significant advancement in ACS risk modelling by integrating environmental and clinical data, offering a comprehensive approach to preventive cardiology. The findings suggest that future predictive models should include environmental data to enhance accuracy and provide more holistic insights into patient health, thereby aiding in more effective clinical decision-making and targeted interventions.
Our future work will includes continuously validating TabulaTime across diverse geographic settings to enhance its generalizability, accounting for variations in air pollution levels, socio-demographics, and healthcare access. Additionally, incorporating personalized risk assessments—factoring in genetic predisposition, lifestyle, and comorbidities—could improve individualized ACS prediction. Beyond air pollution, integrating broader environmental and socioeconomic factors, such as healthcare accessibility, would further refine the model into a comprehensive health risk assessment tool. These advancements will enhance TabulaTime’s impact on clinical decision-making and public health strategies.

### v2

@@ 6.1 Model Performance, Innovations, and Clinical Insights
The robustness of TabulaTime was tested using data from the COVID-19 lockdown period (March 2020–December 2021), when air pollution levels were significantly lower than usual. Despite these different environmental conditions compared to the pre-lockdown training data, the model maintained a high predictive accuracy of 78.8%. This indicates that TabulaTime is adaptable and not solely reliant on the specific pollution patterns present during training. It likely adjusted its predictions, possibly by giving more weight to clinical data when the usual pollution signals were weaker. This ability to handle changing environmental factors highlights the model’s suitability for real-world deployment, such as in health monitoring systems where conditions can vary.
The interpretability features of TabulaTime, derived from attention mechanisms and feature importance analysis, provide valuable clinical insights. The analysis (Figure 8) highlights that established clinical risk factors like Systolic Blood Pressure (SBP), Symptom-to-Admission Time, and Body Mass Index (BMI) are primary drivers in differentiating STEMI from NSTEMI presentations. This aligns with clinical understanding where factors reflecting hemodynamic stability, presentation delay, and metabolic health significantly influence ACS type and severity. Notably, the substantial importance assigned to environmental factors, particularly maximum and average PM10 and average NO, alongside clinical markers like Creatinine and average temperature, underscores a complex interplay between acute environmental triggers, underlying physiological state (e.g., renal function), and the manifestation of specific ACS types. This finding strongly advocates for the integration of environmental monitoring into routine cardiovascular risk assessment and public health strategies.
The temporal analysis (Figure 9) highlights a critical exposure window, with pollution levels 1-3 days prior (peaking at Day 2) showing the highest importance. This strongly indicates that short-term, recent exposure is a significant trigger for acute cardiovascular events, underscoring the clinical and public health value of timely air quality alerts and interventions during high-pollution periods.
The analysis also revealed a slight, though notable, increase in importance for Day 10 compared to Day 9. From a modelling standpoint, in sequential models such as PatchRWKV, earlier time steps in the input sequence (e.g., Day 10) may receive relatively stronger attention weights due to their positional encoding or sequence alignment, potentially influencing feature importance scores. This effect can sometimes exaggerate the influence of positions within fixed-length temporal windows. From a clinical perspective, the finding may reflect a lagged biological response to air pollution exposure. Prior studies have reported delayed effects of environmental pollutants on vascular inflammation, endothelial dysfunction, and thrombotic pathways [37, 67, 68], which may increase vulnerability to ACS several days after exposure. Furthermore, histopathological evidence suggests that plaque erosion—a mechanism responsible for  25% of ACS cases—may initiate approximately 7 days prior to clinical presentation, as evidenced by remodelling of the thrombus overlying the disrupted plaque [69, 70]. It is tempting to speculate that the increased peak around 10 days prior to clinical presentation may represent a plaque erosion population, however this would require further evidence to support such a hypothesis. Smoking is the only known modifiable risk factor for plaque erosion, which could potentially be regarded as the most severe ‘air pollution’ model available.
6.2 Limitations and Future Work
The TabulaTime framework is a flexible and extensible multimodal learning model that integrates both tabular and time-series data for acute coronary syndrome (ACS) subtype prediction. Its modular design enables the incorporation of diverse clinical and environmental inputs, as demonstrated by the seamless addition of external temperature data. TabulaTime outperformed traditional models and effectively captured short-term pollution effects, offering new insights into environmental and physiological contributions to ACS risk.
While TabulaTime has demonstrated strong performance and clinical relevance, there remain several opportunities for further enhancement.
First, the current model was developed using data from a single air quality monitoring station in Salford (UKA00339), which limited the inclusion of some environmental variables. Key meteorological factors such as humidity, atmospheric pressure, rainfall, and wind direction—known to influence pollution dynamics and cardiovascular outcomes—were not available at the time of analysis. Nevertheless, the successful integration of external temperature data illustrates the model’s adaptability, and other climatic variables will be considered in future iterations. Geographic and collective factors such as traffic density and high-resolution socioeconomic status (SES) indicators were also not included due to data unavailability, but they represent valuable additions to improve spatial granularity and context-awareness.
Second, the model was evaluated in a single urban setting. Although Salford represents a socioeconomically diverse population (mean Townsend score  4), broader validation is needed to ensure generalisability across other regions. Encouragingly, the planned expansion to Greater Manchester will support the inclusion of spatially resolved environmental and demographic data, enhancing the model’s robustness and applicability in diverse real-world settings.
Third, although cardiovascular risk scores such as QRISK and the Framingham Risk Score (FRS) are referenced, no direct comparisons were made, as these tools are designed for long-term population-level prediction. In contrast, TabulaTime addresses a specific classification task among patients already diagnosed with ACS. Nevertheless, its architecture offers potential for extension to broader population-level risk prediction in future studies.
Fourth, a major consideration in the current study is the class imbalance between STEMI (n = 123) and NSTEMI (n = 587) cases. Conducting subgroup analyses on this already skewed dataset, particularly within the smaller STEMI group, would result in very small sample sizes per stratum. This introduces a high risk of overfitting, statistical noise, and unreliable effect estimates, which would compromise the validity of any conclusions drawn from such analyses. We believe that performing subgroup analysis under these conditions could be misleading rather than informative. Therefore, we have chosen not to report subgroup-specific performance in the current study, in order to avoid drawing potentially unstable or spurious conclusions. However, we recognise the importance of subgroup insights and are actively planning targeted subgroup analyses in future work, supported by expanded datasets to ensure more robust and meaningful conclusions.
Future work will focus on expanding the model’s geographic scope and input feature set. This includes incorporating additional meteorological variables (e.g., wind direction, pressure, rainfall), as well as spatial and socioeconomic factors such as traffic density and area-level deprivation metrics. The model will be validated across the Greater Manchester region using more comprehensive and spatially resolved datasets. As we expand the cohort to include a larger and more balanced population, we intend to perform a comprehensive investigation of model performance across clinically meaningful subgroups (e.g., by age, cardiovascular history, and risk factor burden), ensuring fairness and generalisability. Such work will be essential for supporting personalised deployment of the TabulaTime model in real-world settings. Furthermore, we explicitly state our intention to apply TabulaTime to multimodal clinical time series datasets, such as those found in MIMIC-IV, eICU, or wearable health platforms. These data sources offer ideal testbeds to demonstrate how TabulaTime can support applications such as early deterioration prediction, dynamic patient risk profiling, and real-time decision support. Finally, as more population-based data become available, TabulaTime may be adapted for primary prevention use cases and directly compared with traditional cardiovascular risk scoring tools for incident ACS prediction.
7 Conclusion
This study introduced TabulaTime, a novel multimodal deep learning framework designed to integrate time-series environmental data (air pollution and climate) with clinical tabular data for enhanced prediction of Acute Coronary Syndrome (ACS) presentation (STEMI vs. NSTEMI). TabulaTime effectively leverages the strengths of both data modalities, offering a more comprehensive approach to cardiovascular risk assessment than traditional clinical-only models. A key innovation is the PatchRWKV module, which provides efficient and accurate automatic feature extraction from time-series data, capturing complex temporal patterns with linear computational complexity.
Our experimental evaluation demonstrated TabulaTime’s superior performance. It significantly outperformed established machine learning models (Random Forest, LightGBM, CatBoost) by 20.5% to 32.2% in accuracy for differentiating ACS subtypes. Crucially, the integration of environmental data improved TabulaTime’s accuracy by 10.1%, highlighting the substantial contribution of air pollution and climate factors to predictive modelling. The PatchRWKV module also excelled independently, surpassing state-of-the-art models in benchmark time-series classification and forecasting tasks. Furthermore, TabulaTime showed robustness, maintaining high accuracy even during the COVID-19 lockdown period despite significant changes in air pollution levels. Feature importance analysis confirmed the relevance of both clinical factors (e.g., Systolic Blood Pressure, Symptom-to-Admission Time, BMI) and environmental variables (e.g., PM10, NO, temperature), particularly emphasizing the impact of short-term pollution exposure in the days preceding an ACS event.
Future work will focus on expanding the model’s geographic scope, incorporating a richer set of environmental and socioeconomic features, and validating its performance across diverse populations and larger, more balanced datasets. Planned subgroup analyses will explore performance variations, and the framework will be applied to other complex clinical time-series datasets (e.g., MIMIC-IV, eICU) and potentially adapted for primary ACS prevention.
In conclusion, TabulaTime represents a significant advancement in integrating environmental health data with clinical risk factors. Its accuracy, efficiency, and interpretability offer considerable potential for improving clinical decision support, informing public health strategies, and advancing personalised cardiovascular medicine.